# FLIP: Towards Fine-grained Alignment between ID-based Models and Pretrained Language Models for CTR Prediction

### v3

@@
Abstract
Click-through rate (CTR) prediction plays as a core function module in various personalized online services. The traditional ID-based models for CTR prediction take as inputs the one-hot encoded ID features of tabular modality, which capture the collaborative signals via feature interaction modeling. But the one-hot encoding discards the semantic information conceived in the original feature texts. Recently, the emergence of Pretrained Language Models (PLMs) has given rise to another paradigm, which takes as inputs the sentences of textual modality obtained by hard prompt templates and adopts PLMs to extract the semantic knowledge. However, PLMs generally tokenize the input text data into subword tokens and ignore field-wise collaborative signals. Therefore, these two lines of research focus on different characteristics of the same input data (i.e., textual and tabular modalities), forming a distinct complementary relationship with each other. In this paper, we propose to conduct Fine-grained feature-level ALignment between ID-based Models and Pretrained Language Models (FLIP) for CTR prediction. We design a novel joint reconstruction pretraining task for both masked language and tabular modeling. Specifically, the masked data of one modality (i.e., tokens or features) has to be recovered with the help of the other modality, which establishes the feature-level interaction and alignment via sufficient mutual information extraction between dual modalities. Moreover, we propose to jointly finetune the ID-based model and PLM for downstream CTR prediction tasks, thus achieving superior performance by combining the advantages of both models. Extensive experiments on three real-world datasets demonstrate that FLIP outperforms SOTA baselines, and is highly compatible for various ID-based models and PLMs. The code is available [^1] https://github.com/justarter/FLIP.
Index Terms:
@@ Abstract
I Introduction
@@
Click-through rate (CTR) prediction is the core component of various personalized online services (e.g., web search [1, 2, 3], recommender systems [4, 5]). It aims to estimate a user’s click probability towards each target item, given a particular context [6, 7, 8]. Recently, the emergence of Pretrained Language Models (PLMs) [9] has facilitated the acquisition of extensive knowledge and enhanced reasoning abilities, hence introducing a novel paradigm for CTR prediction task [10].
@@
The traditional ID-based models for CTR prediction adopt the one-hot encoding to convert the input data into ID features, which we refer to as tabular data modality. An example is shown as follows:
@@
$\underbrace{[0,0,1,\cdots,0]}_{\texttt{UserID=02}}\ \underbrace{[0,1]}_{\texttt{Gender=Male}}\ \underbrace{[0,1,\cdots,0]}_{\texttt{ItemID=01}}\ \underbrace{[0,1,\cdots,0]}_{\texttt{Genre=Action}}$ (1)
@@
Deriving from POLY2 [11] and FM [12], the key idea of these models is to capture the high-order feature interaction patterns across multiple fields with different operators (e.g., product [13, 14], convolution [15, 16], and attention [17, 18]). On the other hand, the development of PLMs has given rise to another model paradigm that utilizes the PLM as a text encoder or recommender directly. The input data is first transformed into textual sentences via hard prompt templates, which could be referred to as textual data modality. In this way, the semantic information among original data is well preserved as natural languages (e.g., feature as text “male” instead of ID code “[0,1]”). Then PLMs [19, 20] encode and understand the textual sentences, and thus turn CTR prediction into either a binary text classification problem [21, 22] or a sequence-to-sequence task [23, 24] for inferring user preference. However, both of the above models possess inherent limitations.
In terms of ID-based models, they utilize the one-hot encoding that discards the semantic information conceived in the original features and fails to capture the semantic correlation between feature descriptions [25, 26]. In addition, ID-based models rely heavily on the user interactions and may not work well in scenarios with sparse interactions [27]. The aforementioned issues can be effectively alleviated by PLMs. PLMs have the ability to extract personalized information from feature descriptions, and use their external knowledge and reasoning capabilities to achieve robust performance in sparse-interaction situations [28, 29].
Furthermore, PLMs also have certain limitations. Since the input data is formulated as textual sentences which would be broken down into subword tokens, language models generally lose sight of field-wise collaborative signals among multiple features [30, 31], which are indispensable in CTR prediction. Additionally, it is hard to distinguish features that exhibits little textual variations in natural language [32, 33] (e.g., in terms of the movie name, “The Room” and “Room” are two very similar movies literally), which hinders language models from capturing subtle similarities and differences between different feature descriptions [34, 35]. Fortunately, the ID-based model can distinguish each different feature and perceive field-wise collaborative signals with the distinctive encoding method and model structure.
@@
Figure 1: Three cross-modal pretraining tasks. Task (a) provides coarse-grained instance-level alignment via contrastive learning, while tasks (b) and (c) achieve fine-grained feature-level alignment through joint reconstruction.
As discussed above, both types of models possess their own advantages and disadvantages. So it is evident to construct a hybrid model that aligns and exploits both collaborative and semantic signals from ID-based models and PLMs, respectively. Meanwhile, in order to effectively solve the above challenges, the alignment of the two modalities needs to be refined to the feature level, so that ID-based models can perceive semantic information corresponding to each feature and PLMs can clearly distinguish features with similar text but different IDs.
However, as shown in Figure 1(a), most existing methods [25, 30] perform the contrastive learning [34] to align the ID-based models and PLMs. The Instance-level Contrastive Learning (ICL) is only coarse-grained and are not sufficient to solve the above feature-level challenges. In addition, ICL can not sufficiently capture the underlying relationships between features. It results in the representation degeneration problem [36, 37, 38], which will degrade the performance. We illustrate the feature embeddings learned by ICL in Figure 2(a), which collapse into a narrow cone and lead to limited expressive capacity. In contrast, our proposed FLIP is able to learn more distributed representations, and therefore better capture the feature diversity as shown in Figure 2(b).
@@
To this end, in this paper, we propose to conduct Fine-grained feature-level ALignment between ID-based Models and Pretrained Language Models (FLIP) for CTR prediction. FLIP is a model-agnostic framework that adopts the common pretrain-finetune scheme [7, 19]. As illustrated in Figure 1(b) and 1(c), for pretraining objectives, we propose to build joint masked tabular and language modeling, where the masked features (i.e., ID or text of some features) of one modality are recovered with the help of another modality. This is motivated by the fact that both tabular and textual data convey almost the same information of the original raw data, but only in different formats. In order to accomplish the masked feature reconstruction, each single model (ID-based model or PLM) is required to seek and exploit the valuable knowledge embedded in the other model that corresponds to the masked features. In this way, the fine-grained feature-level alignment between feature IDs and texts is established.
(a) ICL
(b) Our FLIP
Figure 2: Visualization of feature embeddings on MovieLens-1M dataset. We use SVD [39] to project the feature embedding matrix into 2D data. (a) Instance-level Contrastive Learning (ICL) learns feature embeddings that collapse into a narrow cone, resulting in limited expressive capacity. (b) Our FLIP learns feature embeddings that are more distributed, therefore better capturing the feature diversity.
Then, as for finetuning, the fine-grained alignment pretraining has provided an enhanced modality-aware parameter initialization. Hence, we propose to jointly tune ID-based model and PLM to achieve superior performance. In the experiment, we further explore the effectiveness of finetuning only one of two models.
@@
  • We are the first to highlight the importance of utilizing fine-grained feature-level alignment between ID-based models and PLMs, in order to explore the potential of enhancing the recommendation performance of existing recommender systems.
  • We propose a model-agnostic framework called FLIP, in which the joint masked tabular and language modeling tasks are involved. In these tasks, the masked features of one modality have to be recovered with the help of another modality, thus establishing feature-level interactions and alignments between two modalities.
@@
II Related Work
In this section, we will review previous works which are highly related to ours. We will mainly focus on two aspects, including ID-based models and PLMs for CTR prediction.
II-A ID-based Models for CTR Prediction
Click-through rate (CTR) prediction serves as a core function module in personalized online services, including online advertising, recommender systems, etc [40]. Traditional ID-based CTR models follow the common design paradigm: embedding layer, feature interaction (FI) layer, and prediction layer. These models take as input one-hot features of tabular data modality, and employ various interaction functions to capture collaborative signals among features.
Earlier methods are limited in their ability to model fixed-order or low-order feature interactions. For instance, FM-based approaches [41, 42] employ decomposition parameters to model pairwise interactions. Due to the significance of FI in CTR prediction, numerous studies focus on designing novel structures for the FI layer to capture more informative and complex feature interactions. Wide&Deep [43] combines a wide network and a deep network to achieve the advantages of both. DCN [44] and DCNv2 [14] improve Wide&Deep by replacing the wide part with a more expressive cross network to learn explicit high-order feature interactions. DeepFM [5] combines DNN and FM, and xDeepFM [45] extends DeepFM by using a compressed interaction network (CIN) to capture feature interactions in a vector-wise way. Furthermore, various explicit or implicit interaction operators are designed to improve performance. For instance, PNN [13] uses the productive operator, AFN [46] utilizes the logarithmic operator, AFM [18] and AutoInt [17] use the attention operator.
Although the works above can capture the collaborative knowledge effectively, they disregard the original semantic information of the features and rely heavily on extensive user interactions, which constrains their potential applications and hinders their performance.
II-B PLMs for CTR Prediction
Pretrained Language Models (PLMs), as an emerging technology in the field of Natural Language Processing (NLP), are equipped with a vast number of parameters and undergo their rigorous pretraining process [19, 20]. PLMs demonstrate exceptional success in a wide range of tasks, owing to their extensive knowledge and strong reasoning abilities [47, 48, 49]. Inspired by these achievements, the application of language models for recommender systems has received more and more attention and research [10, 50, 51, 52, 53, 54, 55, 56, 57]. Different from one-hot encoding in ID-based models, this line of research needs to convert the raw data into textual modality, thus retaining the original semantic information of features.
Recent efforts to adapt PLMs for CTR prediction have yielded several breakthroughs. For instance, CTR-BERT [22] leverages a two-tower structure, and feeds user and item information into their respective BERT [19] towers for final CTR prediction. PTab [21] pretrains a BERT model based on masked language modeling using dataset text, and then finetunes this BERT on downstream discriminative tasks. P5 [23] uniformly converts different recommendation tasks into text generation tasks where the backbone model is T5 [58]. M6-Rec [59] utilizes M6 [60] as the backbone for downstream adaptation by representing user behavior data as plain texts and converting the tasks to language generation. Recformer [26] utilizes a bidirectional Longformer [61] to encode items which are formulated as textual key-value attribute pairs. Also, TALLRec [31] and related work [62, 63] utilize decoder-only large language models (e.g., LLaMA [64]) to follow instructions and answer the user preference question that are added after the textual input sentences. Other works also explore the application of language models in different settings and scenarios (e.g., zero-shot [65], few-shot [66], fairness [67], and graph learning [29]).
PLMs can extract personalized text information and harness their knowledge and reasoning capabilities to inform ID-based models, thereby amplifying their recommendation performance. However, it is challenging for PLMs to explicitly capture the field-wise collaborative signals among features and distinguish features with subtle textual differences, resulting in a non-negligible performance decline.
To exploit the advantages of both ID-based models and PLMs, existing state-of-the-art methods, such as CTRL [25] and RLMRec [30], utilize the instance-level contrastive learning (ICL) to align these two models. However, the ICL task is coarse-grained and insufficient to promote the model capacity. It also fails to adequately capture the underlying patterns and relationships among features, resulting in representation degeneration problems.
Therefore, we propose to conduct fine-grained feature-level alignment between ID-based models and PLMs via the joint masked tabular and language modeling pretraining tasks. In the finetuning stage, we propose to tune both models, and thus make full use of the advantages of both textual and tabular data modalities to achieve superior CTR prediction performance.
III Preliminaries
III-A ID-based Models for CTR Prediction
@@
In tradition CTR prediction settings, we model the problem as a binary classification task [5, 13]. The dataset with $N$ instances is presented as $\{(\textbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{0,1\}$ (click or not) is the true label indicating user’s actual click behavior. The input $\textbf{x}_{i}$ is usually categorical tabular data with $F$ different fields, which can be formulated as $\textbf{x}_{i}=[x_{i,1},x_{i,2},\ldots,x_{i,F}]$. The goal of CTR prediction is to estimate the click probability $P(y_{i}=1|\textbf{x}_{i})$ based on input $\textbf{x}_{i}$. Generally, traditional ID-based models adopt the common “Embedding & Feature Interaction” paradigm [17, 45], which can be formulated as the following three layers.
Embedding Layer. The input $\textbf{x}_{i}$ is first transformed into high-dimensional sparse features via one-hot encoding [68]. Then we apply the embedding layer to map the one-hot vectors to dense low-dimensional embedding vectors $\textbf{E}_{i}=[\textbf{e}_{i,1},\textbf{e}_{i,2},\cdots,\textbf{e}_{i,F}]\in\mathbb{R}^{F\times D}$, where $D$ is the embedding size, and each feature $x_{i,f}$ is represented as a fixed-length vector $\textbf{e}_{i,f}\in\mathbb{R}^{D}$.
Feature Interaction (FI) Layer. The FI layer takes as inputs the embedding vectors E, and generates a compact representation $\textbf{v}_{i}$ for each instance $\textbf{x}_{i}$. Different structures are designed to capture the feature interaction patterns with different operators (e.g., MLP [5, 43], product [13, 44, 14], and attention [17, 69]).
Prediction Layer. The prediction layer (usually a linear layer or an MLP module) estimates the click probability $\hat{y_{i}}\in[0,1]$ based on the compact representation $\textbf{v}_{i}$, where $\sigma(x)=1/(1+\text{exp}(-x))$ is the sigmoid function. Finally, the model is trained with the binary cross-entropy (BCE) loss in an end-to-end manner:
$\mathcal{L}_{BCE}(y,\hat{y})=-\frac{1}{N}\sum\nolimits_{i=1}^{N}\left[y_{i}\log\left(\hat{y_{i}}\right)+(1-y_{i})\log\left(1-\hat{y_{i}}\right)\right]$ (2)
III-B PLMs for CTR Prediction
@@
As PLMs have shown remarkable success in a wide range of tasks due to their extensive knowledge and reasoning capabilities [20, 49], researchers now tend to leverage their proficiency in natural language understanding and world knowledge modeling to solve the CTR prediction task [10, 50].
@@
Different from ID-based models, (1) PLMs first transform the input data $\textbf{x}_{i}$ into the textual sentence $\textbf{x}^{text}_{i}$ via hard prompt templates. (2) Then the textual sentence $\textbf{x}^{text}_{i}$ is converted into meaningful lexical tokens through tokenization, which are embedded into a vector space. (3) Next, PLMs utilize their transformer-based neural networks to process the token embeddings obtained from the previous step, generating contextually coherent representation $\textbf{w}_{i}$. (4) Finally, we can either add a randomly initialized classification head (usually MLP) on the representation $\textbf{w}_{i}$ to perform binary classification and predict the click label $y_{i}\in\{0,1\}$ [21, 25], or add a language modeling head to do causal language modeling tasks and predict the likelihood of generating the next keyword (e.g., “yes” or “no”) through a verbalizer [23, 31].
IV Methodology
In this section, we present the FLIP framework, where fine-grained feature-level alignments between PLMs and ID-based models are established to promote the downstream CTR performance.
Figure 3: An illustrative example of modality transformation, which converts tabular data into textual data. We first construct a descriptive sentence for each feature field, and then concatenate them to obtain the final textual data.
IV-A Overview of FLIP
Figure 4: The overall framework of our proposed modality alignment pretraining. The instance-level contrastive learning (ICL) module provides coarse-grained instance-level self-supervisions. The joint reconstruction of masked language/tabular modeling (MLM/MTM) requires to recover the masked feature information with the help of the other modality, thereby enabling fine-grained feature-level alignments.
Our proposed FLIP framework follows the pretrain-finetune scheme, which can be divided into the following three major stages:
  • Modality transformation, as depicted in Figure 3, converts the raw data of tabular modality into textual modality. Note that inputs $\textbf{x}_{i}^{tab}$ and $\textbf{x}_{i}^{text}$ contain almost the same information of raw data, but only in different modalities.
  • Modality alignment pretraining, as shown in Figure 4, explicitly interacts between the dual modalities. Specifically, while the instance-level contrastive learning (ICL) module provides coarse-grained instance-level self-supervisions, the joint reconstruction of masked language/tabular modeling (MLM/MTM) could lead to fine-grained feature-level alignments between two models.
  • Supervised finetuning, as illustrated in Figure 5, further adapts the aligned PLM and ID-based model to the downstream task for final CTR prediction.

### v4

# FLIP: Fine-grained Alignment between ID-based Models and Pretrained Language Models for CTR Prediction

@@
Abstract.
Click-through rate (CTR) prediction plays as a core function module in various personalized online services. The traditional ID-based models for CTR prediction take as inputs the one-hot encoded ID features of tabular modality, which capture the collaborative signals via feature interaction modeling. But the one-hot encoding discards the semantic information included in the textual features. Recently, the emergence of Pretrained Language Models (PLMs) has given rise to another paradigm, which takes as inputs the sentences of textual modality obtained by hard prompt templates and adopts PLMs to extract the semantic knowledge. However, PLMs often face challenges in capturing field-wise collaborative signals and distinguishing features with subtle textual differences. In this paper, to leverage the benefits of both paradigms and meanwhile overcome their limitations, we propose to conduct Fine-grained feature-level ALignment between ID-based Models and Pretrained Language Models (FLIP) for CTR prediction. Unlike most methods that solely rely on global views through instance-level contrastive learning, we design a novel jointly masked tabular/language modeling task to learn fine-grained alignment between tabular IDs and word tokens. Specifically, the masked data of one modality (i.e., IDs and tokens) has to be recovered with the help of the other modality, which establishes the feature-level interaction and alignment via sufficient mutual information extraction between dual modalities. Moreover, we propose to jointly finetune the ID-based model and PLM by adaptively combining the output of both models, thus achieving superior performance in downstream CTR prediction tasks. Extensive experiments on three real-world datasets demonstrate that FLIP outperforms SOTA baselines, and is highly compatible with various ID-based models and PLMs. The code is available [^1] PyTorch version: https://github.com/justarter/FLIP. [^2] MindSpore version: https://github.com/mindspore-lab/models/tree/master/research/huawei-noah/FLIP..
@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688106isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systems
1. Introduction
@@
Click-through rate (CTR) prediction is the core component of various personalized online services (e.g., web search (Lin et al., 2021; Fu et al., 2023; Dai et al., 2021), recommender systems (Xi et al., 2023a; Guo et al., 2017)). It aims to estimate a user’s click probability towards each target item, given a particular context (Zhang et al., 2021; Lin et al., 2023b; Xi et al., 2023b). Recently, the emergence of Pretrained Language Models (PLMs) (Qiu et al., 2020) has facilitated the acquisition of extensive knowledge and enhanced reasoning abilities, hence introducing a novel paradigm for predicting CTR directly through natural language (Lin et al., 2023a).
@@
On the one hand, the traditional ID-based models for CTR prediction adopt the one-hot encoding to convert the input data into ID features, which we refer to as tabular data modality. An example is shown as follows:
@@
(1) $\underbrace{[0,0,1,\cdots,0]}_{\texttt{UserID=02}}\ \underbrace{[0,1]}_{\texttt{Gender=Male}}\ \underbrace{[0,1,\cdots,0]}_{\texttt{ItemID=01}}\ \underbrace{[0,1,\cdots,0]}_{\texttt{Genre=Action}}$
@@
Deriving from POLY2 (Chang et al., 2010) and FM (Rendle, 2010), the key idea of these models is to capture the complex high-order feature interaction patterns across multiple fields by different operators (e.g., product (Qu et al., 2016; Wang et al., 2021), convolution (Xin et al., 2019; Liu et al., 2019b), and attention (Song et al., 2019; Xiao et al., 2017)). On the other hand, the development of PLMs has ushered in another modeling paradigm that utilizes the PLM as a text encoder or recommender directly. The input data is first transformed into textual sentences via hard prompt templates, which could be referred to as textual data modality. In this way, the semantic information among original data is well preserved as natural languages (e.g., gender feature as text “male” instead of ID code “[0,1]”). Then, PLMs (Devlin et al., 2018; Liu et al., 2019a) encode and understand the textual sentences, and thus turn CTR prediction into either a binary text classification problem (Liu et al., 2022; Muhamed et al., 2021) or a sequence-to-sequence task (Geng et al., 2022, 2023). However, both of the above models (i.e., ID-based and PLMs) possess inherent limitations.
@@
Figure 1. Three cross-modal pretraining tasks. Task (a) provides coarse-grained instance-level alignment via contrastive learning, while tasks (b) and (c) achieve fine-grained feature-level alignment through jointly masked modality modeling.
ID-based models utilize the one-hot encoding that discards the semantic information included in the textual features and fails to capture the semantic correlation between feature descriptions (Li et al., 2023a; Lin et al., 2024a). In addition, ID-based models rely heavily on the user interactions and may struggle in scenarios with sparse interactions (Schein et al., 2002). The aforementioned issues can be effectively alleviated by PLMs. PLMs excel in understanding the context and meaning behind textual features, and use their knowledge and reasoning capabilities to achieve robust performance in sparse-interaction situations (Zhao et al., 2023; Wei et al., 2023).
Furthermore, PLMs also have certain limitations. PLMs struggle to understand field-wise collaborative signals because their input data is formulated as textual sentences, which are broken down into subword tokens, thus splitting the field-wise features (Ren et al., 2023; Bao et al., 2023). Additionally, PLMs may fail to recognize subtle differences between different feature descriptions (Radford et al., 2021; Shi et al., 2022) because it is hard to distinguish features that exhibit little textual variations in natural language (Zhang et al., 2023b; Harte et al., 2023) (e.g., in terms of the movie name, “The Room” and “Room” are two very similar movies literally). Fortunately, the ID-based model can perceive field-wise collaborative signals with various model structures and distinguish each different feature with the distinctive ID encodings.
To this end, it is natural to bridge two paradigms to leverage the benefits of both modalities while overcoming their limitations. Moreover, the fine-grained feature-level alignment between tabular IDs and word tokens is critical, enabling ID-based models to perceive semantic information corresponding to each feature ID and allowing PLMs to clearly distinguish features with similar text but different IDs. However, most existing methods (Li et al., 2023a; Ren et al., 2023) align both modalities by instance-level contrastive learning, as shown in Figure 1(a). They only rely on the global view and lack supervision to encourage fine-grained alignment between IDs and tokens, potentially causing representation degeneration problems (Qiu et al., 2022; Gao et al., 2019; Li et al., 2020).
@@
In this paper, we propose to conduct Fine-grained Feature-level ALignment between ID-based Models and Pretrained Language Models (FLIP) for CTR prediction. FLIP is a model-agnostic framework that adopts the common pretrain-finetune scheme (Lin et al., 2023b; Devlin et al., 2018). For pretraining objectives, as illustrated in Figure 1(b) and 1(c), we propose to build jointly masked tabular and language modeling, where the masked features (i.e., IDs or tokens of specific features) of one modality are recovered with the help of another modality. This is motivated by the fact that both tabular and textual data convey almost the same information of the original raw data, but only in different formats. In order to accomplish the masked feature reconstruction, each single model (ID-based model or PLM) is required to seek and exploit the valuable knowledge embedded in the other model that corresponds to the masked features, thereby achieving fine-grained feature-level cross-modal interactions. Then, as for finetuning, we propose a simple yet effective adaptive finetuning approach to combine the predictions from both models for downstream CTR estimation.
@@
  • We highlight the importance of utilizing fine-grained feature-level alignment between tabular IDs and word tokens, in order to explore the potential of enhancing the performance of existing recommender systems.
  • We propose a model-agnostic framework FLIP, where the jointly masked tabular and language modeling tasks are involved. In these tasks, the masked features of one modality have to be recovered with the help of another modality, thus learning fine-grained feature-level cross-modal interactions. FLIP also employs an adaptive finetuning approach to combine the predictions from the ID-based model and PLM for improving performance.
@@
2. Preliminaries
2.1. ID-based Models for CTR Prediction
@@
The traditional CTR prediction is modeled as a binary classification task (Guo et al., 2017; Qu et al., 2016), whose dataset is presented as $\{(\textbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{1,0\}$ is the label indicating user’s actual click behavior, and $\textbf{x}_{i}=[x_{i,1},x_{i,2},\ldots,x_{i,F}]$ is categorical tabular data with $F$ different fields. The goal of CTR prediction is to estimate the click probability $P(y_{i}=1|\textbf{x}_{i})$ based on input $\textbf{x}_{i}$.
Generally, ID-based models adopt the “Embedding & Feature Interaction” paradigm (Song et al., 2019; Lian et al., 2018): (1) the input $\textbf{x}_{i}$ is first transformed into one-hot vectors, which are then mapped to low-dimensional embeddings via an embedding layer. (2) Next, Feature Interaction (FI) Layer is used to process the embeddings, compute complex feature interactions and generate a dense representation $\textbf{v}_{i}$. (3) Finally, the prediction layer (usually a MLP module) estimates the click probability $\hat{y_{i}}\in[0,1]$ based on the dense representation $\textbf{v}_{i}$. The ID-based model is trained with the binary cross-entropy (BCE) loss in an end-to-end manner:
(2) $\mathcal{L}_{BCE}(y,\hat{y})=-\frac{1}{N}\sum\nolimits_{i=1}^{N}\left[y_{i}\log\left(\hat{y_{i}}\right)+(1-y_{i})\log\left(1-\hat{y_{i}}\right)\right]$
2.2. PLMs for CTR Prediction
@@
As PLMs have shown remarkable success in a wide range of tasks due to their extensive knowledge and reasoning capabilities (Liu et al., 2019a; Brown et al., 2020), researchers now tend to leverage their proficiency in natural language understanding and world knowledge modeling to solve the CTR prediction task (Lin et al., 2023a; Liu et al., 2023).
@@
Different from ID-based models, (1) PLMs first transform the input data $\textbf{x}_{i}$ into the textual sentence $\textbf{x}^{text}_{i}$ via hard prompt templates. (2) Then the textual sentence $\textbf{x}^{text}_{i}$ is converted into meaningful lexical tokens through tokenization, which are embedded into a vector space. (3) Next, PLMs utilize their transformer-based neural networks to process the token embeddings obtained from the previous step, generating the contextually coherent representation $\textbf{w}_{i}$. (4) Finally, we can either add a randomly initialized classification head (usually MLP) on the representation $\textbf{w}_{i}$ to perform binary classification and predict the click label $y_{i}\in\{0,1\}$ (Liu et al., 2022; Li et al., 2023a), or add a language modeling head to do causal language modeling tasks and predict the likelihood of generating the next keyword (e.g., “yes” or “no”) through a verbalizer (Geng et al., 2022; Bao et al., 2023).
3. Methodology
3.1. Overview of FLIP
Figure 2. The overall framework of our proposed FLIP.
Figure 2 depicts the architecture of FLIP, which consists of three stages: modality transformation, modality alignment pretraining and adaptive finetuning. Firstly, FLIP transforms the raw data from tabular modality into textual modality. Then, in the modality alignment pretraining, we employ the jointly masked language/tabular modeling task to learn fine-grained modality alignments. Lastly, we propose a simple yet effective adaptive finetuning strategy to further enhance the performance on CTR prediction.